OpenStudy (anonymous):

Which shows 602 - 402 being evaluated using the difference of squares method? (4 points) 602 - 402 = (3600 + 1600)(3600 - 1600) = 10,400,000 602 - 402 = 3600 - 1600 = 2,000 602 - 402 = (60 - 40)2 = 202 = 400 602 - 402 = (60 + 40)(60 - 40) = (100)(20) = 2,000

OpenStudy (anonymous):

i'll give a medal for the first one who ever helps

OpenStudy (welshfella):

a^2 - b^2 = (a + b)(a - b)

OpenStudy (welshfella):

now plug in a = 60 and b = 40 and what do you get?

OpenStudy (anonymous):

? dont have a calculator on me sorry

OpenStudy (welshfella):

you dont need one replace a by 60 and b by 40 and you'll find your answer 60^2 - 40^2 = (60 +_____)(60 -____)

OpenStudy (welshfella):

can you fill in the blanks

OpenStudy (anonymous):

can you walk me through this i am a little confused

OpenStudy (welshfella):

the general identity for the difference of 2 squares is a^2 - b^2 = ( a + b)( a - b) where a and b are 2 numbers if you replace a by 60 and b by 40 you'll be able to work out the answer

OpenStudy (welshfella):

60^2 - 40^2 = (60 +_____)(60 -____) = ?

OpenStudy (anonymous):

(60+60)(60-40)??

OpenStudy (welshfella):

no the second number in first parentheses is b and b = 40

OpenStudy (anonymous):

so the last answer

OpenStudy (anonymous):

D right ?

OpenStudy (welshfella):

60^2 - 40^2 = (60 +__40___)(60 -_40___)

OpenStudy (welshfella):

yes

OpenStudy (welshfella):

a^2 - b^2 = (a + b)(a - b) remember this and you can do any difference of squares problem

OpenStudy (anonymous):

thank you so much for helping me with this you gave me a step by step process that i now understand thanks again

OpenStudy (anonymous):

can you help me on another problem?

OpenStudy (welshfella):

for ex ample 100 - 81 = 10^2 - 9^2 = (10+9)(10-9) = 19*1 = 19

OpenStudy (welshfella):

yes but plz post it separately